# PSCon: Toward Conversational Product Search

### v1

@@ Abstract.
Conversational Product Search (CPS) is confined to simulated conversations due to the lack of real-world CPS datasets that reflect human-like language. Additionally, current conversational datasets are limited to support cross-market and multi-lingual usage. In this paper, we introduce a new CPS data collection protocol and present PSCon, a novel CPS dataset designed to assist product search via human-like conversations. The dataset is constructed using a coached human-to-human data collection protocol and supports two languages and dual markets. Also, the dataset enables thorough exploration of six subtasks of CPS: user intent detection, keyword extraction, system action prediction, question selection, item ranking, and response generation. Furthermore, we also offer an analysis of the dataset and propose a benchmark model on the proposed CPS dataset.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: In Proceedings of XXXXXXX; 978-1-4503-XXXX-X/18/06isbn: ;
@@ 1. Introduction
As for the existing research on Conversational Product Search (CPS), Zhang et al. (2018) presented the first CPS model by simulating conversations over item “aspects” extracted from user reviews. Based on that, Bi et al. (2019) proposed a Conversational Product Search (CPS) model based on the negative feedback and simulated user feedback on the item “aspects.” Similarly, Zou et al. (2022b) simulated conversations based on item “aspects” and proposed a CPS model via representation learning. On the other hand, Zou and Kanoulas (2019) proposed a question-based Bayesian product search model by simulating conversations based on extracted informative terms instead of item “aspects.” The existing work on CPS mainly simulates user conversations while interacting with the product search system. That is, there is a lack of a real CPS dataset driven by human-like language, and thus a public dataset for product search with human-like conversations is of great importance and significance.
@@
To this end, in this paper, we create PSCon, a new dataset for product search with human-like conversations that supports dual markets and two languages. Specifically, we build the dataset by formulating the task of CPS. We define a pipeline for CPS, which includes the following six sub-tasks: (T1) user intent detection, (T2) keyword extraction, (T3) system action prediction, (T4) question selection, (T5) item ranking, and (T6) response generation. First, user intent detection aims to identify the user’s general intent. Second, for each utterance from the user, the system needs to understand its fine-grained information by extracting keywords. Third, the system needs to decide which action to take at each turn, e.g., whether the system should ask a clarifying question or return products to users. Fourth, when the system decides to ask a clarifying question, question selection focuses on selecting high-quality questions to ask the user for preference elicitation. Fifth, when the system decides to return products to users, the system selects products based on conversation history and returns the high-quality products to users. Sixth, response generation incorporates useful information and generates natural language responses to communicate with the user. To collect the CPS dataset, we use a coached human-human data collection protocol to ensure human-like dialogues and avoid algorithm bias (Radlinski et al., 2019), where some participants mimic system roles, i.e., digital shopping assistants, while others act as user roles, i.e., customers, similar to the setting of existing conversational datasets  (Bernard and Balog, 2023). The user role expresses their demands of products and the system role assists them to find their target products through conversations.
@@
To the best of our knowledge, this is the first effort for the CPS dataset with human-like language that is available for dual markets and two languages.
@@ 2.1. Conversational Product Search
Recently, there has been a growing interest regarding CPS in the information retrieval community (Javadi et al., 2023; Montazeralghaem and Allan, 2022; Papenmeier and Topp, 2023; Ye et al., 2024; Li et al., 2025). In the early stage, Zhang et al. (2018) proposed a first CPS model by introducing a unified framework for conversational search and recommendation. This model engages users by posing questions about various aspects derived from user reviews and collects their feedback on these aspect values. Similarly, Bi et al. (2019) simulated conversations and solicited explicit user responses to aspect-value pairs mined from product reviews, with a particular emphasis on the importance of negative feedback in the context of CPS. In contrast to this approach, Zou and Kanoulas (2019) focused on querying users about informative terms, typically entities, extracted from item-related descriptions and reviews. They introduced a sequential Bayesian method that employs cross-user duet training for enhancing CPS. Based on that,  Zou et al. (2020b) conducted an empirical study to measure user willingness and the accuracy of answers provided in question-based product search systems, shedding light on the effectiveness of existing question-based product search approaches. More recently, Zou et al. (2022b) asked clarifying questions (Vedula et al., 2024) on item aspects and proposed a CPS model via representation learning.
@@ 2.2. Conversational Datasets
As for product search and e-commerce, there are only a few conversational datasets available. Fu et al. (2020) built a conversational dataset for e-commerce constructed from Amazon reviews. However, they use simulated users and conversations. Jia et al. (2022) introduced a conversational dataset from the E-commerce domain with user profiles, but they are in Chinese only and involve a single market only. Bernard and Balog (2023) collected a dataset of multi-goal conversations for e-commerce, which is the closest dataset to ours. However, they are in a sole language, i.e., English, and involve a single market only. Also, they only contain less than 20 conversations per category. Different from the above publications, we build a CPS dataset through real human-human conversations, which is on a suitable scale. Moreover, it can support dual markets and two languages, along with a knowledge graph.
@@ 2.3. Conversational Recommender System for e-Commerce
Conversational recommender systems for e-commerce are also associated with product search. For instance, Zhang et al. (2018) propose a unified paradigm for product search and recommendation. Conversational recommender systems utilize human-like natural language to deliver personalized and engaging recommendations through conversational interfaces like chatbots and intelligence assistants (Gao et al., 2021; Jannach et al., 2021). In general, the existing research on conversational recommender systems for e-commerce is usually for anchor-based conversational recommender systems, which are in the form of “system ask–user respond” mode. In this context, the systems ask questions, prompting users to provide feedback, subsequently utilizing this user-provided information to refine their recommendations. They simulate conversations based on the predefined anchors to characterize items, including intent slots (e.g., item aspects and facets) (Zhang et al., 2018), entities (Zou et al., 2020a), and attributes (Zhang et al., 2022).
In contrast to the existing work for conversational recommendation that mainly focuses on preference elicitation and top-k recommendation (Radlinski et al., 2019; Sun and Zhang, 2018), our CPS focuses on item ranking and incorporates user intent detection, keyword extraction, system action prediction, and response generation. That is, they only cover part of aspects of the CPS pipeline.
@@
(a) Search interface.
(b) Chat interface.
Figure 2. The interface of the chat room (the system role).
@@ 3.1. Protocol
    For each round of conversation, the system roles can input a query (i.e., keywords) to search for related results for products. They can also search for products by using search filters provided in the system when necessary.
@@
We value participants’ privacy. All participants are informed that their data is securely encrypted and any identifiable information is removed. We did not collect any data to breach their privacy. We also ask the participants to report any adult content and/or offensive content in the conversations that might raise ethical issues. Our study was approved by the ethics committee of the institute.
@@ 3.2.1. Administration Pages
When entering the system, the new participant needs to register first. When registration, the new participant is assigned a role (i.e., the user role or the system role) randomly, following the random assignment setup of related literature (Harvey and Pointon, 2017; Kelly, 2009; Zou et al., 2022a). After logging in with the assigned role, participants are directed to the instruction page. The instruction page shows detailed instructions regarding the study and what is expected concerning conversational goals. The instructions are designed for each role. That is, the participants of user roles only see their own instructions for customers. In contrast, the participants of system roles only see their own instructions for the system role as a shopping assistant. Following the instructions, some demographic questionnaires are shown to the participants, such as their gender, age, and career field. After that, the participants are redirected to the chat room (see Section 3.2.2 for details). After the conversation in the chat room is done, the participants are redirected to a post-task page. The post-task page shows some post-task questionnaires regarding the experience and satisfaction of participants. In addition, there is a relevance judgment page before the post-task questionnaires. On the relevance judgment page, the user role needs to evaluate the relevance of the products (using like/dislike labels to indicate their assessment of the recommended items) recommended by the system in the conversations.
@@ 3.2.2. Chat Room
The interface for the chat room differs depending on the role of the participant. The interface for the system role is shown in Figure 2, while the interface for the user role is omitted due to its simplicity and limited space. Both the interface for the user role and the system role are divided into two vertical blocks. The left block displays the ongoing conversation while the right block shows the control panel. They both contain an instruction in the top right corner. The interface for the control panel differs between the user role and the system role. For the user roles, the control panel allows them to select an intent and send messages. They can also finish the conversation anytime by clicking the “finish the conversation” button on the bottom. For the system roles, the control panel also allows them to select a system action and send messages. In addition, the system interface includes a search engine. The search engine produces results of products and attributes by keywords. The system’s role is allowed to submit keywords to search results from the search engine. The search engine we used in this paper is based on Amazon product search API. We use Amazon China [^2] https://www.amazon.cn/ when collecting the Chinese dataset, while using Amazon USA [^3] https://www.amazon.com/ when collecting the English dataset. During the conversation, the system allows them to select products from the entire Amazon product pool to send to the user. The selected product list is comprised of curated products with descriptions and pictures. Both the user and the system can urge their chat partner to speed up, which is shown on the upper right of the interface.
@@ 3.4. Participants
Participants in this study were 465 volunteers recruited through email invitations (students and staff of two universities, one in Europe and one in Asia). Of those participants, 63 were in the pilot study and 402 were in the actual study. The pilot study is deployed to make sure the data collection runs smoothly as expected before the actual data collection. Participants were paid around 2.5 pounds on average for each conversation. In the actual study, 196 participants took part in the Chinese dataset collection and 206 participants took part in the English dataset collection. Their demographic data is briefed as follows: (1) Gender: 163 females, 228 males, 11 non-binary; (2) Age: 232 participants were 18–24 years old, 146 were 25–34, and 24 were older than 35; (3) Career field: 145 in computer science, and 257 in others. [^5] We made efforts to recruit a diverse sample (e.g., recruiting participants from various universities, various continents, and various professional backgrounds), response rates and the availability of participants within our target population led to the observed distribution. Also, we have conducted a preliminary analysis of these demographic variations (including age, gender, and career field) on conversation and relevance judgment. We found that there are no obvious and significant differences (biases) in relevance judgment and conversation statistics (e.g., # average conversation turns and # average search keywords) across these demographics.
@@ 3.5. Data Statistics
Besides the conversation, we also collect the information related to the conversation. We collect the conversation annotations including the user intent, keywords, system action, attributes for clarifying questions, and selected products recommended in the conversation to support our subtasks. [^7] The annotations (ground truth in the dataset) are provided by the recruited participants who act as the system roles or user roles. They provide annotations when they chat with each other, for each round of conversation. The example of a collected conversation is shown in Figure 3. Also, the user preference for products (like/dislike) recommended by the system in the conversation is collected. For each search behavior, the search results for products (top 100 products) and related attributes are also collected. In addition, we also collect information on each product including the URL, product image, product title, product description, product reviews, and product metadata such as category, brand, features, product ratings, and related products (also bought, also viewed, bought together, compared, and sponsored products). Given that many existing studies on product search incorporate external knowledge graphs to capture user preference, we construct a knowledge graph by using the product metadata (i.e., products and their attributes) to support the development of CPS models based on knowledge graphs.
@@ 3.6. Data Analysis
We analyze the PSCon dataset with a focus on the distribution of user intents and system actions, which is shown in Figure 4.888The distribution of intent/action is calculated as the percentage of each corresponding intent/action, i.e., the number of utterances with the corresponding intent/action labels divided by the total number of utterances with any intent/action labels. We observe that, in the Chinese dataset, the top intents selected by users are “Interpret” and “Reveal,” while the top system actions are “Recommend” and “Clarify.” This indicates that systems ask clarifying questions or recommend products in most cases, whereas users reveal their product needs and provide feedback to the system in most cases when interacting with the CPS system. We omit the figures of the English dataset, due to a similar trend and limited space.
Moreover, we explore how the number of liked products found by the system is affected by the number of conversation turns, as shown in Figure 5. We obverse that a higher number of turns leads to a significantly higher number of liked products found. The one-way ANOVA test shows significant differences among different number of turns ($p<0.05$). This might be because the number of recommended products increases with the conversation going on. Also, the system captures more information and is closer to the true user preferences, and thus more likely to recommend the products the user likes, with the conversation going on. This is in line with that the performance of CPS models increases with the increase of the number of conversations (Zou et al., 2022b).
@@ 4. Task Design
Keyword Extraction (T2) Conversations usually contain noisy data and more content than necessary. A user utterance may also not be semantically complete due to ellipsis and anaphora (Dalton et al., 2020). It is challenging to feed the user utterances and conversations to the product search model directly. Therefore, it is important to extract keywords from the user utterance for a language understanding task. Given the conversation history $C$ and the current user utterance $U$, this subtask aims to select a sequence of keywords $T_{U}$, that can best describe the user utterance. Formally, it is defined as learning a mapping: ${C,U}\rightarrow T_{U}$.
System Action Prediction (T3) The action stage helps the system to take appropriate actions at the right time, e.g., when to ask a clarifying question to query new information and when to recommend products. Action prediction has been studied in dialogue systems (Peng et al., 2017) and conversational recommender (Lei et al., 2020). However, action prediction has not yet been studied in CPS. Considering previous work on conversational search and conversational recommender (Ren et al., 2021; Lei et al., 2020) while making a connection to product search, we define the system actions: (1) Clarify: ask questions to clarify the user’s product need for preference elicitation; (2) Recommend: give a suggestion, which can be a certain product or a list of products; (3) Inquire: inquire about the response from the user, e.g., ask the user to rephrase its response if it is unclear; (4) No-answer: cannot find related products to meet user’s need; (5) Chitchat: do not fit other labels, e.g., greetings or other utterances unrelated to the user’s product need. This subtask aims to select an optimal system action $a$, that can help the users locate their target items efficiently, given the context (e.g., conversation history $C$, the current user utterance $U$, and returned candidate items $D$). The subtask is formulated as learning a mapping: ${C,U,D}\rightarrow a$.

### v2

# PSCon: Product Search Through Conversations

@@ Abstract.
Conversational Product Search (CPS) systems interact with users via natural language to offer personalized and context-aware product lists. However, most existing research on CPS is limited to simulated conversations, due to the lack of a real CPS dataset driven by human-like language. Moreover, existing conversational datasets for e-commerce are constructed for a particular market or a particular language and thus can not support cross-market and multi-lingual usage. In this paper, we propose a CPS data collection protocol and create a new CPS dataset, called PSCon, which assists product search through conversations with human-like language. The dataset is collected by a coached human-human data collection protocol and is available for dual markets and two languages. By formulating the task of CPS, the dataset allows for comprehensive and in-depth research on six subtasks: user intent detection, keyword extraction, system action prediction, question selection, item ranking, and response generation. Moreover, we present a concise analysis of the dataset and propose the benchmark model on the proposed CPS dataset. Our proposed dataset and model will be helpful for facilitating future research on CPS.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italy.booktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italyccs: Information systems Users and interactive retrievalccs: Information systems Retrieval models and ranking
@@ 1. Introduction
As for the existing research on Conversational Product Search (CPS), Zhang et al. (2018) presented the first CPS model by simulating conversations over item “aspects” extracted from user reviews. Based on that, Bi et al. (2019) proposed a CPS model based on the negative feedback and simulated user feedback on the item “aspects.” Similarly, Zou et al. (2022b) simulated conversations based on item “aspects” and proposed a CPS model via representation learning. On the other hand, Zou and Kanoulas (2019) proposed a question-based Bayesian product search model by simulating conversations based on extracted informative terms instead of item “aspects.” The existing work on CPS mainly simulates user conversations while interacting with the product search system. That is, there is a lack of a real CPS dataset driven by human-like language, and thus a public dataset for product search with human-like conversations is of great importance and significance.
@@
To this end, in this paper, we create PSCon, a new dataset for product search with human-like conversations that supports dual markets and two languages. Specifically, we build the dataset by formulating the task of CPS. We define a pipeline for CPS, which includes the following six sub-tasks: (T1) user intent detection, (T2) keyword extraction, (T3) system action prediction, (T4) question selection, (T5) item ranking, and (T6) response generation. First, user intent detection aims to identify the user’s general intent. Second, for each utterance from the user, the system needs to understand its fine-grained information by extracting keywords. Third, the system needs to decide which action to take at each turn, e.g., whether the system should ask a clarifying question or return products to users. Fourth, when the system decides to ask a clarifying question, question selection focuses on selecting high-quality questions to ask the user for preference elicitation. Fifth, when the system decides to return products to users, the system selects products based on conversation history and returns the high-quality products to users. Sixth, response generation incorporates useful information and generates natural language responses to communicate with the user. To collect the CPS dataset, we use a coached human-human data collection protocol to ensure human-like dialogues and avoid algorithm bias (Radlinski et al., 2019), where some participants mimic system roles, i.e., digital shopping assistants, while others act as user roles, i.e., customers, similar to the setting of existing conversational datasets  (Bernard and Balog, 2023). The user role expresses their demands for products and the system role assists them in finding their target products through conversations.
@@
To the best of our knowledge, this is the first effort for the CPS dataset with human-like language that is available for dual markets and two languages. The dataset and code are available on https://github.com/JieZouIR/PSCon.
@@ 2.1. Conversational Product Search
Recently, there has been a growing interest regarding CPS in the information retrieval community (Javadi et al., 2023; Montazeralghaem and Allan, 2022; Papenmeier and Topp, 2023; Ye et al., 2024; Li et al., 2025). In the early stage, Zhang et al. (2018) proposed a first CPS model by introducing a unified framework for conversational search and recommendation. This model engages users by posing questions about various aspects derived from user reviews and collects their feedback on these aspect values. Similarly, Bi et al. (2019) simulated conversations and solicited explicit user responses to aspect-value pairs mined from product reviews, with a particular emphasis on the importance of negative feedback in the context of CPS. In contrast to this approach, Zou and Kanoulas (2019) focused on querying users about informative terms, typically entities, extracted from item-related descriptions and reviews. They introduced a sequential Bayesian method that employs cross-user duet training for enhancing CPS. Based on that,  Zou et al. (2020b) and  Ma et al. (2024) conducted empirical studies to measure user interactions in question-based product search systems and CPS, shedding light on the effectiveness of existing question-based product search approaches. More recently, Zou et al. (2022b) asked clarifying questions (Vedula et al., 2024) on item aspects and proposed a CPS model via representation learning.
@@ 2.2. Conversational Datasets
As for product search and e-commerce, there are only a few conversational datasets available. Fu et al. (2020) built a conversational dataset for e-commerce constructed from Amazon reviews. However, they use simulated users and conversations. Jia et al. (2022) introduced a conversational dataset from the E-commerce domain with user profiles, but they are in Chinese only and involve a single market only. Bernard and Balog (2023) collected a dataset of multi-goal conversations for e-commerce, which is the closest dataset to ours. However, they are in a single language, i.e., English, and involve a single market only. Also, they only contain less than 20 conversations per category. Different from the above publications, we build a CPS dataset through real human-human conversations, which is on a suitable scale. Moreover, it can support dual markets and two languages, along with a knowledge graph.
@@ 2.3. Conversational Recommender System for e-Commerce
Conversational recommender systems for e-commerce are also associated with product search. For instance, Zhang et al. (2018) propose a unified paradigm for product search and recommendation. Conversational recommender systems utilize human-like natural language to deliver personalized and engaging recommendations through conversational interfaces like chatbots and intelligence assistants (Gao et al., 2021; Jannach et al., 2021). In general, the existing research on conversational recommender systems for e-commerce is usually for anchor-based conversational recommender systems, i.e., they simulate conversations based on the predefined anchors to characterize items, including intent slots (e.g., item aspects and facets) (Zhang et al., 2018), entities (Zou et al., 2020a), and attributes (Zhang et al., 2022). In contrast to the existing work for conversational recommendation that mainly focuses on preference elicitation and top-k recommendation (Radlinski et al., 2019; Sun and Zhang, 2018), our CPS focuses on item ranking and incorporates user intent detection, keyword extraction, system action prediction, and response generation. That is, they only cover part of the aspects of the CPS pipeline.
@@ 3.1. Protocol
    For each round of conversation, the system roles can input a query (i.e., keywords) to search for related results for products. They can also search for products by using the search filters provided in the system when necessary.
@@
We value participants’ privacy. All participants are informed that their data is securely encrypted and any identifiable information is removed. We did not collect any data to breach their privacy. We also ask the participants to report any adult content and/or offensive content in the conversations that might raise ethical issues.
(a) Search interface.
(b) Chat interface.
Figure 2. The interface of the chat room (the system role).
@@ 3.2.1. Administration Pages
When entering the system, the new participant needs to register first. When registration, the new participant is assigned a role (i.e., the user role or the system role) randomly, following the random assignment setup of related literature (Harvey and Pointon, 2017; Kelly, 2009; Zou et al., 2022a). After logging in with the assigned role, participants are directed to the instruction page. The instruction page shows detailed instructions regarding the study and what is expected concerning conversational goals. The instructions are designed for each role. That is, the participants of each role only see their own instructions for the customers or shopping assistants. Following the instructions, some demographic questionnaires are shown to the participants, such as their gender, age, and career field. After that, the participants are redirected to the chat room (see Section 3.2.2 for details). After the conversation in the chat room is done, the participants are redirected to a post-task page. The post-task page shows some post-task questionnaires regarding the experience and satisfaction of participants. In addition, there is a relevance judgment page before the post-task questionnaires. On the relevance judgment page, the user role needs to evaluate the relevance of the products (using like/dislike labels to indicate their assessment of the recommended items) recommended by the system in the conversations.
@@ 3.2.2. Chat Room
The interface for the chat room differs depending on the role of the participant. The interface for the system role is shown in Figure 2, while the interface for the user role is omitted due to its simplicity and limited space. Both the interface for the user role and the system role are divided into two vertical blocks. The left block displays the ongoing conversation while the right block shows the control panel. They both contain an instruction in the top right corner. The interface for the control panel differs between the user role and the system role. For the user roles, the control panel allows them to select an intent and send messages. They can also finish the conversation anytime by clicking the “finish the conversation” button at the bottom. For the system roles, the control panel also allows them to select a system action and send messages. In addition, the system interface includes a search engine. The search engine produces results of products and attributes by keywords. The system role is allowed to submit keywords to search results from the search engine. The search engine we used in this paper is based on the Amazon product search API. [^2] We use Amazon China (https://www.amazon.cn/) when collecting the Chinese dataset, while using Amazon USA (https://www.amazon.com/) when collecting the English dataset. During the conversation, the system allows them to select products from the entire Amazon product pool to send to the user. The selected product list is comprised of curated products with descriptions and pictures. Both the user and the system can urge their chat partner to speed up, which is shown on the upper right of the interface.
@@ 3.4. Participants
Participants in this study were 465 volunteers recruited through email invitations (students and staff of two universities, one in Europe and one in Asia). Of those participants, 63 were in the pilot study and 402 were in the actual study. The pilot study is deployed to make sure the data collection runs smoothly as expected before the actual data collection. Participants were paid around 2.5 pounds on average for each conversation. In the actual study, 196 participants took part in the Chinese dataset collection and 206 participants took part in the English dataset collection.
@@ 3.5. Data Statistics
Besides the conversation, we also collect the information related to the conversation. We collect the conversation annotations, including the user intent, keywords, system action, attributes for clarifying questions, and selected products recommended in the conversation to support our subtasks. [^5] The annotations (ground truth in the dataset) are provided by the recruited participants who act as the system roles or user roles. They provide annotations when they chat with each other, for each round of conversation. The example of a collected conversation is shown in Figure 3. Also, the user preferences for products (like/dislike) recommended by the system in the conversation is collected. For each search behavior, the search results for products (top 50 products) and related attributes are also collected. In addition, we also collect information on each product including the URL, product image, product title, product description, product reviews, and product metadata such as category, brand, features, product ratings, and related products (also bought, also viewed, bought together, compared, and sponsored products). Given that many existing studies on product search incorporate external knowledge graphs to capture user preference, we construct a knowledge graph by using the product metadata (i.e., users, products, like/dislike relationships between users and products; products, product attributes, and relationships between products and corresponding attributes) to support the development of CPS models based on knowledge graphs.
@@ 3.6. Data Analysis
We analyze the PSCon dataset with a focus on the distribution of user intents and system actions, as shown in Figure 4.666The distribution of intent/action is calculated as the percentage of each corresponding intent/action, i.e., the number of utterances with the corresponding intent/action labels divided by the total number of utterances with any intent/action labels. We observe, in the Chinese dataset, the top intents selected by users are “Interpret” and “Reveal,” while the top system actions are “Recommend” and “Clarify.” This indicates that systems ask clarifying questions or recommend products in most cases, whereas users reveal their product needs and provide feedback to the system in most cases when interacting with the CPS system. We omit the figures of the English dataset, due to a similar trend and limited space.
Moreover, we explore how the number of liked products found by the system is affected by the number of conversation turns, as shown in Figure 5. We observe that a higher number of turns leads to a significantly higher number of liked products found. The one-way ANOVA test shows significant differences among different numbers of turns ($p<0.05$). This might be because the number of recommended products increases as the conversation goes on. Also, the system captures more information and is closer to the true user preferences, and thus more likely to recommend the products the user likes, as the conversation goes on. This is in line with that the performance of CPS models increases with the increase of the number of conversations (Zou et al., 2022b).
@@ 4. Task Design
Keyword Extraction (T2) Conversations usually contain noisy data and more content than necessary. A user utterance may also not be semantically complete due to ellipsis and anaphora (Dalton et al., 2020). It is challenging to feed the user’s utterances and conversations to the product search model directly. Therefore, it is important to extract keywords from the user utterance for a language understanding task. Given the conversation history $C$ and the current user utterance $U$, this subtask aims to select a sequence of keywords $T_{U}$, that can best describe the user utterance. Formally, it is defined as learning a mapping: ${C,U}\rightarrow T_{U}$.
System Action Prediction (T3) The action stage helps the system to take appropriate actions at the right time, e.g., when to ask a clarifying question to query new information and when to recommend products. Considering previous work on conversational search and conversational recommender (Ren et al., 2021; Lei et al., 2020; Zou et al., 2024) while making a connection to product search, we define the system actions: (1) Clarify: ask questions to clarify the user’s product need for preference elicitation; (2) Recommend: give a suggestion, which can be a certain product or a list of products; (3) Inquire: inquire about the response from the user, e.g., ask the user to rephrase its response if it is unclear; (4) No-answer: cannot find related products to meet user’s need; (5) Chitchat: do not fit other labels, e.g., greetings or other utterances unrelated to the user’s product need. This subtask aims to select an optimal system action $a$, that can help the users locate their target items efficiently, given the context (e.g., conversation history $C$, the current user utterance $U$, and returned candidate items $D$). The subtask is formulated as learning a mapping: ${C,U,D}\rightarrow a$.